# SPIRAL: Self-Play on Zero-Sum Games Incentivizes Reasoning via Multi-Agent Multi-Turn Reinforcement Learning

### v2

@@ 6.3 RQ3: Do different games develop specialized reasoning skills?
      TicTacToe Specialist 56.6% 24.4% 30.5% 31.4% 30.8% 30.7% 34.1%
      Kuhn Poker Specialist 31.0% 48.5% 28.7% 24.9% 39.1% 33.9% 34.4%
      Simple Negotiation Specialist 27.7% 16.8% 39.1% 12.3% 32.0% 34.8% 27.1%
Multi-Game Model 54.3% 53.9% 33.2% 51.4% 39.3% 37.2% 44.9%
The power of multi-game training becomes evident in performance on novel games. On Liar’s Dice, individual specialists struggle, achieving only 24.9% (Kuhn Poker) and 12.3% (Simple Negotiation) against Gemini-2.0-Flash-Lite. However, the multi-game model achieves 51.4%, demonstrating emergent capabilities beyond any single training game. This pattern of multi-game superiority holds across the out-of-distribution games, with the multi-game model achieving the highest average performance (44.9%) compared to all specialists.
Multi-Game Training Enhances Even Strong Reasoning Models.
To test whether game-based skill development is limited to weaker models, we apply multi-game SPIRAL training to DeepSeek-R1-Distill-Qwen-7B, a model that already achieves 59.7% average on reasoning benchmarks (see Sec. C.2 for comprehensive results across all training configurations):
Table 6: Multi-game SPIRAL training improves reasoning capabilities across model scales. Even strong reasoning models benefit from game-based skill development.
Model MATH500 AIME’24 AIME’25 OlympiadBench AMC-23 Minerva Math GPQA MMLU-Pro Average
Qwen3-4B-Base 65.8 10.0 3.3 33.3 50.0 24.3 30.6 47.2 33.1
+ SPIRAL (Multi-Game) 74.2 20.0 16.7 39.4 55.0 38.6 36.9 57.7 42.3
DeepSeek-Distill-Qwen-7B 90.8 46.7 36.7 56.9 92.5 48.2 48.6 57.1 59.7
+ SPIRAL (Multi-Game) 93.0 43.3 46.7 57.9 92.5 51.1 49.6 58.9 61.7
Multi-game SPIRAL improves Qwen3-4B by 9.2 percentage points (33.1% to 42.3%) and enhances the already-strong DeepSeek-Distill-Qwen-7B model by 2.0 points (59.7% to 61.7%). This demonstrates that game-based reasoning skills provide value even for models that already excel at mathematical reasoning, suggesting that games teach complementary cognitive abilities not captured by traditional training.
Finding 3: Different games develop specialized skills that work together Each game builds distinct abilities that transfer to similar out-of-distribution games. Multi-game training combines these specialized skills, achieving superior performance on novel games that individual specialists struggle with. This synergistic training improves reasoning benchmarks even for already-strong models, demonstrating that games teach complementary cognitive skills.
6.4 RQ4: Is Role-Conditioned Advantage Estimation (RAE) essential for stable self-play training?
We ablate our variance reduction technique to determine whether RAE is necessary to stabilize training and prevent thinking collapse in zero-sum self-play with shared parameters. Figure 6 reveals that without proper variance reduction, self-play training catastrophically fails:
1. Thinking Collapse: The most striking failure occurs in response length (bottom right). Without RAE, reasoning traces plummet from 3,500 to near-zero characters after 200 steps. Models quite literally stop thinking, generating minimal outputs like <think></think><answer>
boxed{bet}</answer>
. This collapse coincides with gradient norms approaching zero, indicating the model has converged to a degenerate policy.
2. Performance Degradation: Without RAE, the model learns solely to play the game while abandoning CoTs, leading to generalization failure. As a result, the math reasoning performance crashes from 35% to 12% at around step 150 (a 66% relative decrease), and the general reasoning drops from 44% to 40%.
3. Training Instability: Policy gradient norms without RAE exhibit higher initial values with spikes, then collapses to near-zero after step 200. In contrast, RAE maintains stable gradient norms of around 0.1 throughout training, enabling continuous learning.
This variance reduction is crucial for maintaining stable reasoning. High-variance gradients cause models to oscillate between extreme strategies, eventually converging to degenerate short responses that minimize variance but abandon reasoning entirely.
Figure 6: Training dynamics comparing REINFORCE with RAE (orange) versus vanilla REINFORCE (gray). RAE maintains stable performance across all metrics while vanilla REINFORCE suffers catastrophic thinking collapse. Top row: Performance on game playing, math reasoning, and general reasoning benchmarks. Bottom row: Policy gradient norm shows training instability without RAE; response length reveals thinking collapse where models stop generating reasoning traces.
Finding 4: RAE prevents models from abandoning reasoning during self-play Without RAE, models began to truncate their reasoning processes after 200 steps, generating empty reasoning traces like <think></think>. This thinking collapse causes math performance to crash while gradients become unstable then drop to zero. RAE maintains stable gradients throughout training, allowing models to keep generating substantive reasoning. Self-play alone is not enough; proper variance reduction is essential for maintaining reasoning in competitive games.
@@ 7 Conclusion
Summary of Contributions. We introduced SPIRAL, enabling language models to develop reasoning capabilities through competitive self-play without human-curated data. Our technical contributions include a fully online multi-turn MARL system for LLMs and Role-conditioned Advantage Estimation (RAE), which prevents thinking collapse in zero-sum games. Our empirical results show that training on Kuhn Poker alone improves mathematical reasoning by 8.7% average and Minerva Math by 18.1%, surpassing models trained on 25,000 expert demonstrations. Different games develop distinct transferable skills, and multi-game training achieves superior performance on both reasoning benchmarks and novel out-of-distribution games, demonstrating synergistic benefits from diverse game training.
Limitations. While eliminating human-curated problems, SPIRAL still requires designed game environments. Our experiments use simple games (TicTacToe, Kuhn Poker, Simple Negotiation); scaling to complex environments remains unexplored. The computational requirements are substantial (8 H100 GPUs for 25 hours per experiment). Performance plateaus after extended training, and our evaluation focuses on academic benchmarks rather than real-world reasoning tasks requiring common sense or ethical judgment.
Future Directions. This work opens several avenues: expanding to cooperative games, incorporating partial observability, and designing games targeting specific reasoning weaknesses. Understanding why certain games develop particular skills could enable principled environment design. We envision ecosystems of self-improving agents generating increasingly sophisticated challenges, creating autonomous reasoning development beyond human supervision.
Final Thoughts. SPIRAL demonstrates that simple games can unlock complex reasoning without domain-specific data. By harnessing competitive pressure, we create systems that discover their own curricula and continuously improve. The transfer from gameplay to mathematics could suggest that intelligence emerges not from sophisticated supervision but could from environmental challenges that force models to think. This paradigm shift points toward AI systems that autonomously push reasoning boundaries and continuously evolve through self-play.
Appendix
This appendix provides comprehensive details supporting our main findings. App. A shows how existing SFT and RLVR paradigms adapt to our turn-level MDP formulation. App. B contains implementation specifics including hyperparameters, infrastructure details, and game environment observations referenced in Section 5.2. App. C presents extended results tables and detailed pattern evolution analysis that support our case study in Section 6.1. App. D describes the complete methodology for our reasoning pattern analysis, detailing how we discovered and quantified transfer mechanisms. Finally, App. E provides comprehensive descriptions of for all game environments used in our experiments, including both training and evaluation games.
Appendix A Training Paradigms in Turn-Level MDPs
This section shows how existing training paradigms adapt to our turn-level MDP formulation introduced in Section .
A.1 Supervised Fine-Tuning (SFT) in Turn-Level MDPs
@@
A.2 Reinforcement Learning with Verifiable Rewards (RLVR) in Turn-Level MDPs
@@
In single-turn settings without subsequent interactions, RLVR reduces to a contextual bandit problem (Langford and Zhang, 2007). Recent works on mathematical (Shao et al., 2024; DeepSeek Team, 2024) and code reasoning (Zhu et al., 2024; Xin et al., 2024) show that even this simplified bandit-style RLVR can unlock sophisticated reasoning. However, these approaches still require human-curated problem sets $\mathcal{D}_{\text{RLVR}}$, which SPIRAL eliminates through self-play.
Appendix B Experimental Setup Details
@@
B.1 Game Environment Observations
The language models receive structured text observations from each game environment. Fig. 7 shows example observations from our three training games: TicTacToe, Kuhn Poker, and Simple Negotiation. These observations serve as the input prompts $s_{t}$ at each turn, providing complete game state information in natural language format.
@@ B.1 Game Environment Observations
Similarly, Fig. 8 presents observations from five evaluation environments used to test out-of-distribution generalization. These games were never seen during training, allowing us to assess whether learned skills transfer to novel game mechanics.
@@
B.2 Hyperparameter Configuration
Table 7 presents the complete hyperparameter settings used across all experiments. These configurations were selected through preliminary experiments to balance training stability and computational efficiency. Please see our open-source codebase for a complete and reproducible experiment example.
@@ B.2 Hyperparameter Configuration
Maximum response length $4096$ tokens
@@
These hyperparameters remain fixed across all game environments and model scales to ensure fair comparison. The distributed training infrastructure utilizes 8 H100 GPUs, with parallel actors generating game trajectories while a centralized learner performs synchronous policy updates.
Appendix C Additional Results and Analysis
C.1 Detailed Evolution of Case-by-Case Analysis
To understand how reasoning patterns develop during training, we tracked the evolution of case-by-case analysis across checkpoints. Table 8 shows a concrete example from Minerva Math Problem 135, illustrating how models progressively develop structured reasoning.
Table 8: Evolution of Case-by-Case Analysis in Minerva Problem 135
@@ C.1 Detailed Evolution of Case-by-Case Analysis
[…]
@@ 1. Understanding the Context: BBN at $\sim$20 minutes
 2. Nearest Power of 10: $1010^{1}=$ 10
[…]
@@ 1. Understanding the Context: Recombination at $\sim$377,000 years
 2. Nearest Power of 10: $00010^{5}=100,$ 000
[…]
@@ 1. Understand Context: Current universe age
C.2 Comprehensive Benchmark Results
Table 9 presents extended results showing SPIRAL’s performance across different training configurations and base models.
Table 9: SPIRAL training improves reasoning benchmarks for different base models
Model MATH500 AIME’24 AIME’25 OlympiadBench AMC-23 Minerva Math GPQA MMLU-Pro Average

### v3

@@ 4.1 Understanding Why SPIRAL Works
     TicTacToe Specialist 56.6% 24.4% 30.5% 28.1% 97.6% 79.9% 52.9%
     Kuhn Poker Specialist 31.0% 48.5% 28.7% 27.7% 98.8% 81.6% 52.7%
     Simple Negotiation Specialist 27.7% 16.8% 39.1% 26.4% 98.6% 82.8% 48.6%
Multi-Game Model 54.3% 53.9% 33.2% 31.6% 99.8% 84.0% 59.5%
Figure 6: Training dynamics comparing REINFORCE with RAE (orange) versus vanilla REINFORCE (gray). RAE maintains stable performance across all metrics while vanilla REINFORCE suffers catastrophic thinking collapse. Left: Response length reveals thinking collapse where models stop generating reasoning traces; Right: Performance on general reasoning benchmarks.
[…]
@@ 5 Conclusion
We introduced SPIRAL, enabling language models to develop reasoning capabilities through competitive self-play without human-curated data. Our technical contributions include a fully online multi-turn MARL system for LLMs and Role-conditioned Advantage Estimation (RAE), which prevents thinking collapse in zero-sum games. Empirically, multi-game SPIRAL training improves reasoning benchmarks by up to 10.5% across diverse model architectures, surpassing supervised fine-tuning on 25,000 expert game trajectories. Different games develop distinct transferable skills (spatial, probabilistic, strategic) that combine synergistically. Analysis reveals that competitive gameplay forces discovery of reasoning patterns (case-by-case analysis, expected value calculation, pattern recognition) that transfer to academic domains.
SPIRAL demonstrates that simple games can unlock complex reasoning without domain-specific data. Future work could expand to cooperative games, incorporate partial observability, and design games targeting specific reasoning weaknesses. Understanding game-skill mappings could enable principled environment design for autonomous reasoning development.
Acknowledgements
We thank Xidong Feng and Runji Lin for their helpful discussions and support throughout this project. We thank John Schulman for insightful feedback and encouragement. We would like to thank Thinking Machine Lab and Modal Lab for providing compute credits that supported our experiments. We also thank the TextArena team for maintaining the game environments used in this work. This research was supported by the Cooperative AI Foundation, the UW-Amazon Science Gift Hub, Sony Research Award, UW-Tsukuba Amazon NVIDIA Cross Pacific AI Initiative (XPAI), the Microsoft Accelerate Foundation Models Research Program, Character.AI, DoorDash, and the Schmidt AI2050 Fellows program. This material is based upon work supported by the Defense Advanced Research Projects Agency and the Air Force Research Laboratory, contract number(s): FA8650-23-C-7316. Any opinions, findings and conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of AFRL or DARPA.
Reproducibility statement
We provided the experiment code at https://github.com/spiral-rl/spiral. We have also provided the training settings in the experiment section (§4) and Appendix D. The experiments are run with an 8 H100 GPU cluster, and all calls for proprietary LLMs are via the official API or OpenRouter [^3] https://openrouter.ai/.
Appendix A Large Language Model Usage
We used large language models (LLMs) only for language refinement tasks, including grammar checking, phrasing adjustments, and enhancing readability. We also used Deep Research (OpenAI, 2025) to assist with related work search. Besides these, all scientific ideas, experiments, analyses, and results are the sole contributions of the authors.
Appendix B Limitations
Our study, while promising, has several limitations that offer avenues for future research.
Reliance on Designed Game Environments
A core limitation is the dependency on engineered game environments. Although SPIRAL eliminates the need for human-curated problem datasets, it shifts the dependency to well-designed games. The games used in our experiments, such as the Tictactoe and Khun Poker, are relatively simple and feature dense rewards. It is an open question how well this approach scales to more complex, open-ended environments with sparse rewards, such as Minecraft or realistic robotics simulations. The design of the game environment itself may implicitly encode biases or heuristics that influence the agent’s learned reasoning strategies, potentially limiting their generality.
Computational Cost and Scalability
The computational requirements for training are substantial. Each experimental run demanded 8 H100 GPUs for approximately 25 hours, which may be prohibitive for many research groups. Furthermore, we observed that performance gains began to plateau after extended training periods. This suggests that simply scaling up the training duration with the current framework may yield diminishing returns, and more efficient algorithms or architectural improvements are necessary for further progress.
Evaluation and Transferability
Our evaluation, while comprehensive, has two key constraints:
  • Focus on Academic Benchmarks: We primarily assessed reasoning on established academic benchmarks like MATH and GPQA. These benchmarks are excellent for measuring formal and scientific reasoning but do not capture the full spectrum of self-play.
  • Zero-Shot Evaluation: The strict zero-shot evaluation setting tests for direct transfer but may not fully reveal the model’s potential. Fine-tuning on a small set of target domain examples could potentially unlock significantly better performance, a possibility not explored in this work.
Potential for Reward Hacking
Like many reinforcement learning systems, SPIRAL is susceptible to reward hacking. An agent might discover policies that maximize the in-game score without learning the intended underlying reasoning skill. For instance, it could exploit a bug in the game physics or find a repetitive, degenerate strategy that succeeds for a narrow set of problems. While we did not observe significant instances of this, it remains a risk, especially in more complex environments where robust reward shaping is challenging.
Appendix C Preliminaries
This section provides the mathematical foundations and formal definitions underlying the SPIRAL framework, including turn-level MDPs, two-player zero-sum Markov games, and how existing training paradigms adapt to these formulations.
C.1 Turn-level Markov Decision processes (MDPs).
Language model training traditionally formulates generation as a token-level MDP (Bellman, 1957; Rafailov et al., 2024) where each action is a single token from vocabulary $\mathcal{V}$. For multi-turn reasoning and game-playing, we instead adopt a turn-level MDP formulation $\mathcal{M}=(\mathcal{S},\mathcal{A},T,r,\gamma)$. Here, states $\mathcal{S}$ represent complete contexts (e.g., game configurations, problem states, or conversation histories), actions $\mathcal{A}$ are complete responses (containing many tokens), the transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ determines state dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ provides immediate rewards, and $\gamma\in[0,1]$ is the discount factor. The return is defined as the discounted sum of rewards: $R(\tau)=\sum_{t=0}^{T}\gamma^{t}r_{t}$.
where $c_{t}$ externalizes reasoning and $a_{t}\in\mathcal{A}$ is the executable action. (See App.˜C for how existing SFT and RLVR paradigms adapt to turn-level MDPs.)
C.2 Two-Player Zero-Sum Markov Games.
We extend the single-agent MDP to competitive settings with a two-player zero-sum Markov game (Littman, 1994) $\mathcal{G}=(\mathcal{S},\mathcal{A}_{0},\mathcal{A}_{1},T,r,\gamma)$, where $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are the action spaces for player 0 and player 1 respectively. The zero-sum property requires:
C.3 Supervised Fine-Tuning (SFT) in Turn-Level MDPs
@@
C.4 Reinforcement Learning with Verifiable Rewards (RLVR) in Turn-Level MDPs
@@
In single-turn settings without subsequent interactions, RLVR reduces to a contextual bandit problem (Langford & Zhang, 2007). Recent works on mathematical (Shao et al., 2024; DeepSeek Team, 2024) and code reasoning (Zhu et al., 2024; Xin et al., 2024) show that even this simplified bandit-style RLVR can unlock sophisticated reasoning. However, these approaches still require human-curated problem sets $\mathcal{D}_{\text{RLVR}}$, which SPIRAL eliminates through self-play.
Appendix D Experimental Setup Details
@@
D.1 Game Environment Observations
The language models receive structured text observations from each game environment. Fig.˜7 shows example observations from our three training games: TicTacToe, Kuhn Poker, and Simple Negotiation. These observations serve as the input prompts $s_{t}$ at each turn, providing complete game state information in natural language format.
@@ D.1 Game Environment Observations
Similarly, Fig.˜8 presents observations from five evaluation environments used to test out-of-distribution generalization. These games were never seen during training, allowing us to assess whether learned skills transfer to novel game mechanics.
@@
D.2 Hyperparameter Configuration
Table˜6 presents the complete hyperparameter settings used across all experiments. These configurations were selected through preliminary experiments to balance training stability and computational efficiency. Please see our open-source codebase for a complete and reproducible experiment example.
@@ D.2 Hyperparameter Configuration
Maximum response length $8192$ tokens
@@
These hyperparameters remain fixed across all game environments and model scales to ensure fair comparison. The distributed training infrastructure utilizes 8 H100 GPUs, with parallel actors generating game trajectories while a centralized learner performs synchronous policy updates. On this hardware setup, the compute budget per experiment was approximately 25 hours for Qwen3-4B, 28 hours for Qwen3-8B, and 16 hours for both Llama3.1-8B-Instruct and Octothinker-8B. For SFT, we use Qwen3-32B (Yang et al., 2025) models to self-play on the targeted games, generating 25k winning trajectories and crafting them as the SFT dataset.
D.3 Evaluation Settings
To investigate whether the reasoning abilities developed through gameplay could transfer to non-game contexts, we evaluate our models on a suite of established benchmarks. All evaluations on these benchmarks are conducted in a zero-shot setting [^4] Except for the base model, for which we provide few-shot examples that follow the Qwen3 Report settings. to determine if game-induced reasoning could be successfully transferred to general problem-solving. We use a sampling temperature of $0.6$ and top-p of $0.95$ for all evaluations.
Math Benchmarks.
For mathematical reasoning, we use MATH500 (Hendrycks et al., 2021), OlympiadBench (He et al., 2024), Minerva Math (Lewkowycz et al., 2022), AIME24, AIME25 (MAA, a), and AMC23 (MAA, b) datasets, which cover a wide range of topics including algebra, geometry, and competitive mathematics. Following the settings in Zhou et al. (2025), we report avg@32 for AIME24, AIME25 and AMC23; and pass@1 for other math benchmarks.
General Reasoning Benchmarks.
For general reasoning, we utilize GPQA-Diamond (Rein et al., 2024), which consists of graduate-level science questions, and MMLU-Pro (Wang et al., 2024), a benchmark for multidisciplinary knowledge.
Appendix E Additional Results and Analysis
E.1 Detailed Evolution of Case-by-Case Analysis
To understand how reasoning patterns develop during training, we tracked the evolution of case-by-case analysis across checkpoints. Table˜7 shows a concrete example from Minerva Math Problem 135, illustrating how models progressively develop structured reasoning.
Table 7: Case-by-Case Analysis in Minerva Problem 135
@@ E.1 Detailed Evolution of Case-by-Case Analysis
<output> The age when helium formed is approximately 10 minutes… The age when hydrogen formed is 377,000 years… The current age is 13.8 Gyr. Rounding to nearest power of 10, we get: $\boxed{14}$ Gyr </output>
@@ 1. Understanding the Context: BBN at $\sim$20 minutes
 2. Nearest Power of 10: $10^{1}=10$
 3. Final Answer: $\boxed{1}$ minute
@@ 1. Understanding the Context: Recombination at $\sim$377,000 years
 2. Nearest Power of 10: $10^{5}=100,000$
 3. Final Answer: $\boxed{400000}$ years
@@ 1. Understand Context: Current universe age
E.2 Extended Analysis of RAE Ablation Study
Our ablation study of Role-conditioned Advantage Estimation reveals additional insights beyond those presented in Section 4 Figure 9 shows three additional metrics that further demonstrate RAE’s critical role in stable self-play training.
Figure 9: Extended training dynamics comparing REINFORCE with RAE versus vanilla REINFORCE (continued from Figure 6). Left: Game win rates showing RAE achieves faster initial learning compared to vanilla REINFORCE. Middle: Math reasoning performance crashes from 35% to 12% without RAE. Right: Policy gradient norms exhibit instability then collapse to near-zero without RAE, while RAE maintains stable gradients around 0.1.
Game Performance. The left panel shows that REINFORCE with RAE learns significantly faster, rapidly reaching 35% win rate while vanilla REINFORCE exhibits slower initial learning. RAE’s superior learning efficiency demonstrates the benefits of role-conditioned advantage estimation for accelerating policy improvement.
Math Reasoning Collapse. The middle panel reveals the most dramatic failure: math reasoning performance without RAE crashes from 35% to 12% at around step 150 (a 66% relative decrease). This collapse occurs precisely when models stop generating reasoning traces, confirming that thinking collapse directly causes reasoning failure.
Gradient Stability. The right panel shows policy gradient norms, revealing the underlying optimization dynamics. Without RAE, gradients exhibit high variance with erratic spikes before collapsing to near-zero after step 200, indicating convergence to a degenerate policy. RAE maintains stable gradient norms around 0.1 throughout training, enabling continuous improvement.
These additional results reinforce our main finding: self-play alone is insufficient for stable reasoning development. Proper variance reduction through role-specific baselines is essential to prevent models from converging to degenerate policies that abandon reasoning in favor of minimal outputs.
E.3 Game Generalization
We conducted additional experiments on higher-complexity variants of our training games (5x5 Tic-Tac-Toe, 5-Card Kuhn Poker, and 8-Resource Negotiation) to assess scalability.
Table 8: Generalization performance on increased complexity environments. SPIRAL retains significantly higher performance on out-of-distribution (OOD) tasks compared to SFT.
Environment Setting Qwen3-4B Base SFT SPIRAL (Ours)
Tic-Tac-Toe $\times3$ (Train) 17.5 46.9 54.3
$\times5$ (OOD) 5.5 17.8 27.3
Kuhn Poker 3-Card (Train) 21.5 43.6 53.9
5-Card (OOD) 21.9 28.6 50.1
Simple Negotiation 5 Resources (Train) 15.6 26.7 33.2
8 Resources (OOD) 5.5 8.7 31.0
Average (Train) 18.2 39.1 47.1
Average (OOD) 11.0 18.4 36.1
Table 8 shows that although SFT has comparable improvement over the base model in in-domain environments (averaging $39.1\%$ vs. $18.2\%$), it struggles to adapt to OOD cases with increased complexity. The SFT model’s performance drops to only $18.4\%$. In contrast, SPIRAL demonstrates better generalizability. It not only outperforms SFT in the training settings (averaging $47.1\%$) but maintains robust performance in the more complex OOD environments, achieving an average of $36.1\%$. This suggests that SPIRAL is more capable of handling increased complexity.
E.4 Comprehensive Benchmark Results
Table˜9 presents extended results showing SPIRAL’s performance across different training configurations and base models.
Table 9: SPIRAL training improves reasoning benchmarks for different base models. We include an additional baseline of SFT on 52k trajectories to demonstrate that simply scaling supervised data does not match the gains from self-play.
Model MATH500 AIME’24 AIME’25 OlympiadBench AMC-23 Minerva Math GPQA-D MMLU-Pro Average